### v1

@@ 2 Preliminaries
For any mechanism $M$ and $x\in\mathbb{R}^{\times}$,
$h_{M}(x)=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M)}[(1-xe^{-Z})_{+}]$
where $(t)_{+}:=\max\{0,t\}$.
Proof.
$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-xQ(E)$
$=\sup_{E\subseteq\Omega}\int_{E}1-x/\frac{dP}{dQ}(\omega)\,dP(\omega)$
$=\int_{\Omega}\left(1-x/\frac{dP}{dQ}(\omega)\right)_{+}\,dP(\omega)$
$=\int_{\Omega}(1-xe^{-\log(\ell(\omega))})_{+}\,dP(\omega)$
For convenience, we will sometimes write $h_{L}$ to denote the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x):=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$. Moreover, like PLDs, hockey-stick curves also capture adaptive composition in a natural way.
@@
Proof.
By Proposition˜5 and the law of total expectation, it follows that
Our last characterization of privacy is given the Type I/Type II error tradeoff curve, also known as the $f$-DP framework of privacy [1].
@@
Let $(P,Q)$ be any distributions on $\Omega$. For any hypothesis test $\phi:\Omega\to\{P,Q\}$ for distinguishing $P$ and $Q$, we call $\alpha_{\phi}:=\mathbb{P}_{\omega\sim P}[\phi(\omega)=Q]$ its Type I error and $\beta_{\phi}:=\mathbb{P}_{\omega\sim Q}[\phi(\omega)=P]$ its Type II error. The tradeoff curve for $(P,Q)$ is defined as
@@
For convenience, we sometimes write the tradeoff curve of a mechanism $M$ with respect to the fixed dataset pair $(D_{1},D_{2})$ as $T_{\alpha}(M):=T_{\alpha}(M(D_{1})\parallel M(D_{2}))$ or sometimes as $\tau_{M}(\alpha):=T_{\alpha}(M)$ where appropriate. In general, we can characterize valid tradeoff curves as follows.
Proposition 7 ([1] Proposition 2.2).
A curve $\tau:[0,1]\to[0,1]$ is a tradeoff curve for some testing problem $(P,Q)$ exactly when $\tau$ is continuous, convex, decreasing, and $\tau(\alpha)\leq 1-\alpha$ for all $\alpha\in[0,1]$.
Like hockey-stick curves, it is enough to understand the PLD of a testing problem in order to obtain its tradeoff curve. The following formula can be derived by the well-known Neyman–Pearson lemma.
@@
Let $(P,Q)$ be a pair of distributions with tradeoff curve $\tau(\alpha):=T_{\alpha}(P\parallel Q)$ and let $F$ denote the CDF of $\operatorname{PLD}(P\parallel Q)$. Then
$\tau(\alpha)=\mathbb{E}_{Z\sim\operatorname{PLD}(P\parallel Q)}[1(F(Z)>\alpha)\cdot e^{-Z}]$
It is common to represent a pair of distributions by either their hockey-stick curve or their tradeoff function. We show that there is a natural link between these representations via inversion and convex conjugacy. For a tradeoff function $\tau$ of a testing problem $(P,Q)$, we write $\tau^{-1}(\beta):=\inf\{\alpha\in[0,1]:\tau(\alpha)\leq\beta\}$ for its inverse, namely the tradeoff curve for the testing problem $(Q,P)$. For a convex function $g$, we denote its convex conjugate by $g^{*}(y):=\sup_{x\in\mathbb{R}}xy-g(x)$. Note that we can extend tradeoff functions to the real-line by setting them to $\infty$ outside their support $[0,1]$. Note that a special case of the following formula appears in [1] (Prop 2.12) for symmetric tradeoff functions but this generalization is new to the best of our knowledge.
@@
Proof.
First, let $L$ denote the PLD of $(P,Q)$ and let $F$ be its CDF. By Proposition˜8,
$\tau(\alpha)=\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot e^{-Z}].$
Therefore, by Proposition˜5
$h(x)$ $=\mathbb{E}_{Z\sim L}[(1-xe^{-Z})_{+}]$
$=\mathbb{E}_{Z\sim L}[1(Z>\log(x))\cdot\underbrace{(1-xe^{-Z})}_{>0\iff Z>\log(x)}]$
$=\sup_{\varepsilon\in\mathbb{R}}\mathbb{E}_{Z\sim L}[1(Z>\varepsilon)\cdot(1-xe^{-Z})]$
$=\sup_{\alpha\in[0,1]}\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot(1-xe^{-Z})]$
$=\sup_{\alpha\in[0,1]}1-\alpha-x\tau(\alpha)$
$=\sup_{\alpha^{\prime}\in[0,1]}1-\tau^{-1}(\alpha^{\prime})-x\alpha^{\prime}$ ($\alpha^{\prime}=\tau^{-1}(\alpha)$)
$=1+(\tau^{-1})^{*}(-x).$
@@
Let $(P,Q)$ and $(P^{\prime},Q^{\prime})$ be pairs of distributions with tradeoff functions $\tau$ and $\tau^{\prime}$ respectively as well as hockey-stick curves $h$ and $h^{\prime}$ respectively. The following are equivalent
  1. (i)
    $(P,Q)\preceq(P^{\prime},Q^{\prime})$
  2. (ii)
    $\tau\succeq\tau^{\prime}$
  3. (iii)
    $h\preceq h^{\prime}$
The equivalence of (i) and (ii) is given by a celebrated theorem of Blackwell [Bla51, Theorem 10]. As for the equivalence of (ii) and (iii), this now follows by applying Fenchel–Moreau duality to Proposition˜9 and recalling that convex conjugation is an order-reversing operation.
@@
For any non-empty family of PLDs $\mathcal{L}$ dominated by at least one PLD, there exists a unique PLD $\sup\mathcal{L}$ that dominates $\mathcal{L}$ but is dominated by every upper bound for $\mathcal{L}$. In this case,
@@
This result follows immediately from Proposition˜10 and from noticing that the properties characterizing hockey-stick curves (see Proposition˜4) are all closed under pointwise suprema. Part (ii) follows from noticing that all of the properties of tradeoff curves (see Proposition˜7) are also closed under the lower convex envelope operation.
We rely on the supremum extensively in order to reason about natural privacy filters. In particular, the former characterization says that the supremum of PLDs is given by a pointwise supremum in the space of hockey-stick curves, which will be particularly useful for showing our main result. As far as we are aware, the supremum property of PLDs has not actually been documented in the literature.
@@ 3 Query-Restricted Adversaries
In this section we develop a general theory of the privacy characteristics of adversarial adaptive mechanism when the queries the adversary may issue are restricted. This will lead directly to a theory of natural filters as we will see in §4.
@@
Let $k$ be a positive integer. We define a privacy rule of length $k$ as a map $\Gamma$ that takes a sequence of PLDs $(L_{1},\dots,L_{k^{\prime}})$ ($0\leq k^{\prime}<k$) and outputs some collection of PLDs including Id. We define a $\Gamma$-adversary as an adaptively composed mechanism $M=M_{1}\otimes\dots\otimes M_{k}$ such that
$L_{i}:=\operatorname{PLD}(M_{i}(\cdot;y_{1},\dots,y_{i-1}))\in\Gamma(L_{1},\dots,L_{i-1})$
@@
It turns out that the maximum privacy loss for an adversary respecting a rule $\Gamma$ occurs in some sense when the adversary first plays a move and then adversarially chooses a optimal strategy tuned to ensure that the overall strategy is optimal at a particular given point. Note that by $\lambda$ we mean the empty sequence.
@@
Let $\Gamma$ be a rule of length $k$. If $k=1$, then $\operatorname{PLD}(\Gamma)=\sup{\Gamma(\lambda)}$ and, for $k>1$, we have
$\operatorname{PLD}(\Gamma)=\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}$
where $\Gamma_{L_{0}}(L_{1},\dots,L_{k^{\prime}}):=\Gamma(L_{0},L_{1},\dots L_{k^{\prime}})$ denotes the rule of length of k1k-1 that fixes the first move played against $\Gamma$ to $L_{0}$.
@@
$\operatorname{PLD}(\Gamma)\preceq\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}.$
@@
To that end, consider any $\Gamma$-adversary $M$. Decompose $M$ as the adaptive composition of $M_{1}(\cdot)$ with some $M^{\prime}(\cdot;y_{1})$ and let $L:=\operatorname{PLD}(M_{1})\in\Gamma(\lambda)$. Then, clearly, for every fixed $y_{1}$, $M^{\prime}(\cdot;y_{1})$ is a $\Gamma_{L}$-adversary and, in particular, $\operatorname{PLD}(M^{\prime}(\cdot;y_{1}))\preceq\operatorname{PLD}(\Gamma_{L})$. By Proposition˜3, we have that $\operatorname{PLD}(M)\preceq L\otimes\operatorname{PLD}(\Gamma_{L})$ and the desired inequality follows by taking suprema.
@@
More surprising is the reverse inequality, which we prove constructively by passing to hockey-stick curves. Let $L\in\Gamma(\lambda)$, let $\kappa\in\mathbb{R}^{\times}$, and let $\varepsilon>0$. Choose any mechanism $M_{1}$ that has PLD $L$ and let $\ell:=\frac{dM_{1}(x_{1})}{dM_{1}(x_{2})}$ denote the corresponding likelihood function. Moreover, for any $\kappa\in\mathbb{R}^{\times}$ we can choose by Proposition˜11, a $\Gamma_{L}$-adversary $M^{\prime}_{\kappa}$ such that
$H_{\kappa}(M^{\prime}_{\kappa})\geq\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}H_{\kappa}(M^{\prime\prime})-\varepsilon.$
Now consider the $\Gamma$-adversary
$M(x):=(Y_{1},Y^{\prime}),Y_{1}\sim M_{1}(x),Y^{\prime}\sim M^{\prime}_{\kappa/\ell(Y_{1})}(x).$
By Proposition˜6, we get
$h_{M}(\kappa)$ $=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[h_{M^{\prime}_{\kappa/\ell(Y_{1})}}(\kappa/\ell(Y_{1}))]$
$\geq\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}}(\kappa/\ell(Y_{1}))-\varepsilon]$
$=(h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}})(\kappa)-\varepsilon.$
Again, taking suprema, we get that
$\sup_{M\in\operatorname{Adv}(\Gamma)}h_{M}\succeq h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}},$

### v2

@@ 2 Preliminaries
For any pair $(P,Q)$ with likelihood $\ell$ and $x\in\mathbb{R}^{\times}$,
$H_{x}(P\parallel Q)=\mathop{\mathbb{E}}_{Y\sim P}[(1-x/\ell(Y))_{+}]=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[(1-xe^{-Z})_{+}].$
For convenience, we will sometimes write $h_{L}$ to denote the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x):=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$. Moreover, like PLDs, hockey-stick curves also capture adaptive composition in a natural way. The proof is also in Appendix˜A.
@@
Our last characterization of privacy is given the Type I/Type II error tradeoff curve, also known as the $f$-DP framework of privacy Dong et al. (2019).
@@
Let $(P,Q)$ be distributions on $\Omega$. For a randomized hypothesis test $\phi:\Omega\to\Delta(\{P,Q\})$, we call $\alpha_{\phi}:=\mathbb{P}_{\omega\sim P,\mu\sim\phi(\omega)}[\mu=Q]$ its Type I error and $\beta_{\phi}:=\mathbb{P}_{\omega\sim Q,\mu\sim\phi(\omega)}[\mu=P]$ its Type II error. The tradeoff curve for $(P,Q)$ is defined as
@@
In general, we can characterize valid tradeoff curves as follows.
Proposition 7 (Dong et al. (2019) Proposition 2.2).
A curve $\tau:[0,1]\to[0,1]$ is the tradeoff curve for some $(P,Q)$ iff $\tau$ is continuous, convex, decreasing, and $\tau(\alpha)\leq 1-\alpha$ for all $\alpha\in[0,1]$.
As with hockey-stick curves, the PLD determines the tradeoff curve. The well-known Neyman–Pearson lemma states that the optimal testing rule thresholds the PLD. For convenience, given $z^{*}\in\overline{\mathbb{R}}$, we call $\psi:\overline{\mathbb{R}}\to[0,1]$ a $z^{*}$-threshold when $\psi(z)=0$ for $z<z^{*}$ and $\psi(z)=1$ for $z>z^{*}$.
@@
Let $(P,Q)$ be a pair of distributions and let $F$ be the CDF of $\operatorname{PLD}(P\parallel Q)$. Then, for every $\alpha\in[0,1]$, setting $z_{\alpha}:=\min\{z\in\overline{\mathbb{R}}:F(z)\geq\alpha\}$, there is a $z_{\alpha}$-threshold $\psi_{\alpha}$ such that
$\tau(\alpha)=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[\psi_{\alpha}(Z)e^{-Z}]$
and such that $\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[\psi_{\alpha}(Z)]=1-\alpha$. Moreover, if $\alpha=F(z^{*})$, then, for $Z\sim\operatorname{PLD}(P\parallel Q)$, $\psi_{\alpha}(Z)=1(Z>z^{*})$ almost surely.
For convenience, we denote the tradeoff curve corresponding to a PLD $L$ as $\tau_{L}$. It is common to represent the privacy characteristics of a mechanism by either its hockey-stick curve or its tradeoff curve. We show that there is a natural link between these representations via inversion and convex conjugacy. For a tradeoff function $\tau$ of a testing problem $(P,Q)$, we write $\tau^{-1}(\beta):=\inf\{\alpha\in[0,1]:\tau(\alpha)\leq\beta\}$ for its inverse, namely the tradeoff curve for the testing problem $(Q,P)$. For a convex function $g$, we denote its convex conjugate by $g^{*}(y):=\sup_{x\in\mathbb{R}}xy-g(x)$. Note that we can extend tradeoff functions to the real-line by setting them to $\infty$ outside their support $[0,1]$. Note that a special case of the following formula appears in Dong et al. (2019) (Prop 2.12) for symmetric tradeoff functions but this generalization is new to the best of our knowledge. The proof is in Appendix˜A
@@
Let $(P,Q)$ and $(P^{\prime},Q^{\prime})$ be pairs of distributions with tradeoff functions $\tau$ and $\tau^{\prime}$ respectively as well as hockey-stick curves $h$ and $h^{\prime}$ respectively. Then
$(P,Q)\preceq(P^{\prime},Q^{\prime})\iff\tau\succeq\tau^{\prime}\iff h\preceq h^{\prime}.$
The equivalence of (i) and (ii) is given by a celebrated theorem of Blackwell (Blackwell, 1951, Theorem 10). As for the equivalence of (ii) and (iii), this now follows by applying Fenchel–Moreau duality to Proposition˜9 and recalling that convex conjugation is an order-reversing operation.
@@
For any non-empty family of PLDs $\mathcal{L}$, there exists a unique PLD $\sup\mathcal{L}$ that dominates $\mathcal{L}$ but is dominated by every upper bound for $\mathcal{L}$. In this case,
@@
This result follows immediately from Proposition˜10 and from noticing that the properties characterizing hockey-stick curves (see Proposition˜4) are all closed under pointwise suprema. Part (ii) follows from noticing that all of the properties of tradeoff curves (see Proposition˜7) are also closed under the lower convex envelope operation. Notice that it may be the case that $\sup{\mathcal{L}}=\infty$. Also notice that we may construct infima of privacy losses in an analogous manner, namely by taking lower convex envelopes of hockey-stick curves or by taking pointwise suprema of tradeoff curves. We will not require infima for our work but we do rely on the supremum extensively in order to reason about natural privacy filters. As far as we are aware, the completeness property of PLDs has not actually been documented in the literature.
@@ 3 Query-Restricted Adversaries
In this section we develop a general theory of the privacy characteristics of an adversarial adaptive mechanism when the queries the adversary may issue are restricted. This will lead directly to a theory of natural filters as we will see in §5. For simplicity, we will assume throughout that an adversary may always “pass” its turn by issuing a query with no privacy loss. We will consider adversaries that issue finitely many queries as well as adversaries that issue a countable sequence of queries. We will abuse notation somewhat by writing such things as $M_{1}\otimes\dots\otimes M_{\infty}$ to denote the countable extension of the prefixes $M_{1}\otimes\dots\otimes M_{t}$ as $t\to\infty$.
@@
Let $k\in\mathbb{Z}^{+}\cup\{\infty\}$. We define a privacy rule of length $k$ as a map $\Gamma$ that takes a sequence of PLDs $(L_{1},\dots,L_{t})$ ($0\leq t<k$) and outputs some collection of PLDs including Id. We define a $\Gamma$-adversary as an adaptively composed mechanism $M=M_{1}\otimes\dots\otimes M_{k}$ such that
$L_{t}:=\operatorname{PLD}(M_{t}(\cdot;y_{1},\dots,y_{t-1}))\in\Gamma(L_{1},\dots,L_{t-1})$
@@
Crucially, notice that that no single adversary fully realizes the privacy cost of the rule $\Gamma$. Recalling the $f$-DP interpretation of the supremum (see Proposition˜11), the privacy cost of the rule $\Gamma$ is in fact realized by a family of adversaries that minimize the false negative rate for every given confidence level. As the following core lemma reveals, it turns out that the maximum privacy cost of the rule $\Gamma$ for a given confidence level occurs when the adversary first plays the best move for this confidence level, observes the outcome, then depending on the likelihood of having observed this outcome recursively chooses its remaining strategy tuned to maximize the overall privacy cost. Note that by $\lambda$ we mean the empty sequence.
@@
Let $\Gamma$ be a rule of length $k\in\mathbb{Z}^{+}\cup\{\infty\}$. If $k=1$, then $\operatorname{PLD}(\Gamma)=\sup{\Gamma(\lambda)}$ and, for $k>1$, we have
$\operatorname{PLD}(\Gamma)=\sup\{L\oplus\operatorname{PLD}(\Gamma|_{L}):L\in\Gamma(\lambda)\}$
where $\Gamma|_{L_{0}}(L_{1},\dots,L_{t}):=\Gamma(L_{0},L_{1},\dots L_{t})$ denotes the rule of length of k1k-1 that fixes the first move played against $\Gamma$ to $L_{0}$.
We prove the result by passing to hockey-stick curves because the pointwise supremum is easier to analyze than the lower convex envelope. We will essentially show that a worst-case adversary first plays an optimal move with corresponding PLD $L$, observes the outcome $Y_{1}$, and adversarially chooses a followup in $\operatorname{Adv}(\Gamma|_{L})$ to ensure that the overall process “traces out” the hockey-stick curve of $L\oplus\operatorname{Adv}(\Gamma|_{L})$.
@@
$\operatorname{PLD}(\Gamma)\preceq\sup\{L\oplus\operatorname{PLD}(\Gamma|_{L}):L\in\Gamma(\lambda)\}.$
@@
To that end, consider any $\Gamma$-adversary $M=M_{1}\otimes M_{2}$ and let $L:=\operatorname{PLD}(M_{1})\in\Gamma(\lambda)$. Then, clearly, for every fixed $y_{1}$, $M_{2}(\cdot;y_{1})$ is a $\Gamma|_{L}$-adversary and, in particular, $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(\Gamma|_{L})$. By Proposition˜3, $\operatorname{PLD}(M)\preceq L\oplus\operatorname{PLD}(\Gamma|_{L})$ and thus the desired inequality follows by taking suprema.
@@
More surprising is the reverse inequality, which we prove constructively using hockey-stick curves. Let $L\in\Gamma(\lambda)$, $x\in\mathbb{R}^{\times}$, and $\gamma>0$. Choose any mechanism $M_{1}$ with PLD $L$ and let $\ell_{1}:=\frac{dM_{1}(D_{1})}{dM_{1}(D_{2})}$ denote its likelihood function. Moreover, for any $x^{\prime}\in\mathbb{R}^{\times}$ we can choose by Proposition˜11, a $\Gamma|_{L}$-adversary $M^{x^{\prime}}$ such that $h_{M^{x^{\prime}}}(x^{\prime})\geq h_{M^{\star}}(x^{\prime})-\gamma$ where $\operatorname{PLD}(M^{\star})=\operatorname{PLD}(\Gamma|_{L})$. Now consider the $\Gamma$-adversary $M:=M_{1}\otimes M_{2}$ where $M_{2}(D;y_{1}):=M^{x/\ell_{1}(y_{1})}(D)$. By Proposition˜6, we get
$h_{M}(x)$ $=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{x/\ell_{1}(Y_{1})}}(x/\ell_{1}(Y_{1}))]$
$\geq\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M^{\star}}(x/\ell_{1}(Y_{1}))-\gamma]$
$=h_{M_{1}\otimes M^{\star}}(x)-\gamma.$